# OFER: Occluded Face Expression Reconstruction

### v1

@@ 4.2 Identity Ranking Network (IdRank)
Our goal is to obtain a set of 3D faces with diverse expressions that match the input occluded image. While the expressions can vary, the underlying geometric structure of the face must remain unique and consistent for each identity. Thus, given the variability among the generated shape samples obtained from IdGen, the next essential step is to select the best representative shape for the identity from these samples. However, we observed that the shape coefficients display much less variability than expression coefficients (see Sec. 4.3), as images typically contain more cues for the expression-independent components of the face. Given the occlusion ambiguity and limited shape variability, multiple samples may approximate the ground truth and thus be considered equally optimal. For this reason, we chose to rank the samples, as ranking improves robustness by evaluating top options instead of relying on a potentially suboptimal single choice.
@@
To achieve this, we introduce IdRank, $R$, a model specifically designed to identify the sample that best preserves the facial structure of the input image.
Sample Quantity.
The goal of the ranking network is to select, among a pool of candidate shapes, the one that is closest in distance to the (unknown) ground-truth shape. In the extreme case, given an infinite number of samples, one of them being an exact match to ground truth would provide zero reconstruction error. Thus, the probability of IdRank selecting a shape closer to the ground truth relies on having an ample pool of options to begin with. Producing a large number of samples is computationally infeasible; hence, it is important to find a good trade-off between computational complexity and the probability of encountering a ’good’ shape within the samples. For this paper, we empirically choose a sample size of $100$ (see supplementary).
Preprocessing.
Efficiently selecting the most accurate sample also requires a network design that excludes irrelevant information. Because the goal is to match the facial features of the input image, information from the back of the head is irrelevant. Including it in vertex displacement error metric computation may lead to incorrect sample selection. Therefore, we use only the frontal face vertices, $x_{ssf}$, as input to this network.
Ground truth.
In order to establish the ground truth in IdRank, for each training image, we preprocess and extract the front face vertices from the IdGen-generated samples, $X_{ssf}=\{x_{ssf_{1}},x_{ssf_{2}},\ldots,x_{ssf_{n}}\}$, and the corresponding ground truth face training data, referred to as $x_{sgf}$. The ground truth for the ranking network is then obtained as a sorted ‘ordered’ set by the following process:
$x_{sgf_{i}}=x_{sgf}-x_{ssf_{i}}$ (2)
$X_{sgf}=\{x_{sgf_{1}},x_{sgf_{2}},\ldots,x_{sgf_{N}}\}{}$ (3)
$d_{x_{sgf_{i}}}=\frac{1}{m}\sum_{j=0}^{m}|v_{j}(x_{sgf})-v_{j}(x_{ssf_{i}})|{}$ (4)
[…]
such that
$\quad\operatorname{rank}(x_{sgf_{1}}^{\prime})\geq\operatorname{rank}(x_{sgf_{2}}^{\prime})\geq\ldots>\operatorname{rank}(x_{sgf_{N}}^{\prime})$ (6)
Here, $X_{sgf}^{\prime}$ represents the same set as $X_{sgf}$ but ordered in ascending order based on the vertex displacement error, such that $d_{x_{sgf_{1}}^{\prime}}\leq d_{x_{sgf_{2}}^{\prime}}\leq\ldots\leq d_{x_{sgf_{N}}^{\prime}}$. We define the elements of $X_{sgf}$ and $X_{sgf}^{\prime}$ such that $x_{sgf_{i}}\in X_{sgf}$ corresponds to to $x_{sgf_{p}}^{\prime}\in X_{sgf}^{\prime}$.
Input.
The input to the ranking network is a set of mean-residual pairs, where mean ($X_{\text{mean}}$) of all generated samples is concatenated with the residual ($x_{\text{res}_{i}}$) of each sample forming $\hat{x}_{i}=(X_{\text{mean}},x_{\text{res}_{i}})$. This creates the final input set for the network, denoted as $\hat{X}=\{\hat{x}_{1},\hat{x}_{2},\ldots\hat{x}_{N}\}$. This approach is motivated by the fact that removing the mean reduces redundancy, enhancing feature representation through residual values [32]. Our ablation study on type of input to the network (see Tab. 7) supports this.
Training.
We train $R$ to match the ranking order of $X_{ssf}$ to the ranked order of $X_{sgf}^{\prime}$. During each training iteration, for each image, a new set of $X_{ssf}$ shape samples is generated by IdGen to capture most of the feature variances. The network is conditioned on a 1024-dimensional feature vector $\boldsymbol{C_{\textit{R}}}=(c_{a}\|c_{f})$, formed by concatenating features from the ArcFace ($c_{a}$) and FaRL ($c_{f}$) encoders. The network predicts a score, $p_{i}$, for each mean-residual pair sample $\hat{x}_{i}$. The set of all sample scores for an image is represented as $P=\{p_{1},p_{2},\ldots p_{N}\}$. The objective is to sort $P$ and optimize a loss based on ranking order. We choose a list-wise ranking method (ref. Sec. 3) due to its effectiveness [7].
Since sorting and ranking are non-differentiable operations, we reformulate the sort-rank problem into a probability distribution alignment problem. We achieve this by converting the ground truth vertex error values ( Eq. 4) of $\hat{X}_{sgf}$ and $P$ into probability distributions $g$ and $h$. With this reformulation, the network optimizes the parameter $\theta_{R}$ with gradient back-propagation to IdGen disabled, aiming to minimize the softmax loss $L_{R}$ between the ground truth distribution $g$ and the predicted distribution $h$.
$R(\hat{x}_{i}|C_{R};\boldsymbol{\theta_{R}})=p_{i}$ (7)
$h(\hat{X})=\frac{\mathrm{exp}(p_{i})}{\sum_{j=1}^{N}\mathrm{exp}(p_{j})}{}$ (8)
$g(X_{sgf})=\frac{\mathrm{exp}(d_{x_{sgf_{i}}})}{\sum_{j=1}^{N}\mathrm{exp}(d_{x_{sgf_{i}}})}{}$ (9)
$L_{R}(g,h;\boldsymbol{\theta_{R}})=-\sum_{i=1}^{N}g(X_{sgf_{i}})\log(h(\hat{X}_{i})){}$ (10)
@@
During inference, we select the sample with the highest probability score for a given occluded input image as the optimal sample, denoted as $S_{R}$. While multiple samples may have equal probability scores, their ordering in such cases would be random rather than favoring a specific sample. Therefore, we choose the first-ranked sample from this list of equally high-probability samples.
Figure 5: Identity Ranking Network. Top row: Starting from unit Gaussian noise $\epsilon\in\mathbb{R}^{300}$, we sample N shape coefficients from trained IdGen for given image. FLAME reconstructs the entire head shape from these coefficients (ref. Sec. 3). We preprocess and extract frontal face vertices (ref. Sec. 4.2) from it obtaining the set $X_{ssf}=\{x_{ssf_{1}},x_{ssf_{2}},\ldots,x_{ssf_{N}}\}$. Bottom row: The mean of $X_{ssf}(X_{\text{mean}})$ concatenated with the residual of each sample $x_{ssf_{i}}(x_{\text{res}_{i}})$ is given as input to the Ranking Network $R$. The network, conditioned on the image, optimizes its parameters to match ranking orders of $X_{ssf}$ to that of the ground truth (Eq. 5).
@@ 4.3 Expression Generative Network (ExpGen)
Figure 6: Architecture of Expression Generative Network. During training, the input image is encoded into a 512-dimensional embedding using the ArcFace [11] and another 512-dimensional embedding using FaRL [53]. The concatenated embedding is used as a condition for the 3-layer 1D U-Net Diffusion Network. The ground truth 50-dimensional 3DMM expression coefficients of the image is passed as input to the network.
The expression network, ExpGen, is trained in parallel with IdGen and IdRank. Its architecture and sample generation process closely follow that of IdGen, utilizing the conditional DDPM framework to generate expression coefficients. We train ExpGen using the first 50 PCA components of the FLAME expression coefficients, conditioning on a 1024-dimensional feature vector, similar to IdRank, expressed as $\boldsymbol{C_{\textit{E}}}=(c_{a}||c_{f})$.
During sampling, the network initializes from unit Gaussian noise $\varepsilon\in\mathbb{R}^{50}$ at $t=T$ and iteratively denoises it to approach the FLAME expression parameters ($e_{0}=\hat{E}\in\mathbb{R}^{50}$). The network optimizes the parameters $\theta_{E}$ by minimizing the loss function $L_{\varepsilon}$ as defined below:
$L_{\varepsilon}(\boldsymbol{\theta_{\textit{E}}})=\mathbb{E}_{t}\Bigl{[}\mathbb{E}_{e_{0},\varepsilon}[|\hat{\varepsilon}_{t}-\varepsilon_{t}|]\Bigr{]}$ (11)
4.4 Final Loss and Reconstruction
All the network parameters of OFER are optimized by minimizing the overall loss $L$ defined as:
$L=L_{\varepsilon}(\boldsymbol{\theta_{\textit{S}}})+L_{\varepsilon}(\boldsymbol{\theta_{\textit{R}}})+L_{\varepsilon}(\boldsymbol{\theta_{\textit{E}}})$ (12)
The final reconstruction of diverse expressive faces for a given occluded image is obtained by passing the shape coefficients of the optimal sample $S_{R}$, selected by IdRank and each of the $N$ sampled expression coefficients obtained from ExpGen through the FLAME model. Qualitative results of these final reconstructions are shown in Fig. 1.
@@ 5 Experiments and Evaluations
In this section, we introduce the datasets used in our method and provide details of the experimental configuration, evaluation metrics, comparison with baseline methods. Finally, we present qualitative and quantitative results for the three networks detailed in Sec. 4.
@@ 5.1 Datasets
Training.
We used the Stirling [23], FaceWarehouse [8], LYHM [9], and Florence [3] datasets to train IdGen. These datasets consists of 2D-3D ground truth pairs of images and corresponding registered FLAME shape coefficients. The same four datasets were used to train IdRank. For ExpGen, we utilized the FaMoS (Facial Motion across Subjects) [6] dataset, which offers paired image and FLAME 3D meshes. The decoupled shape and expression coefficients for each pair were provided by the FaMos authors. Detailed information on the number of identities and images from each dataset used in our experiments is available in the supplementary materials.
5.2 Identity Generative Network Evaluation
Experimental Setup.
We use the four datasets detailed in Sec. 5.1 for training the network and the NoW benchmark [45] for validation. The network is trained for 400 epochs with a batch size of 128 on an RTX 8000 GPU. The entire training process spans one GPU day.
@@ 5.2 Identity Generative Network Evaluation
We compare our approach with state-of-the-art occlusion-based reconstruction methods [14, 31]. Additionally, we randomly sample FLAME shape coefficients from their parametric space to serve as a baseline benchmark. Our evaluation metrics include the mean, median, and standard deviation of the mean square error (MSE) as provided by the NoW benchmark [45]. We also provide a comprehensive comparison with other neutral face reconstruction methods [55, 52, 50, 22, 45, 16, 24] in the supplement.
@@ Results.
Method Average error of samples
Med $\downarrow$ Mean $\downarrow$ std $\downarrow$
FLAME(Baseline) Without ranking 1.79 2.25 1.90
FOCUS [31] 1.05 1.31 1.14
FOCUS (MP) [31] 1.03 1.29 1.12
Diverse3DFace[14] 1.41 1.78 1.52
OFER (ours) 1.02 1.25 1.04
OFER (sample selected by ranking) 0.98 1.21 1.01
Table 1: Comparison of non-metrical evaluation in NOW validation benchmark. FLAME (baseline) is created by sampling random FLAME shape coefficients, which serves as our baseline. We evaluate against occlusion reconstruction methods using the median (Med), mean, and standard deviation (std) of the non-metrical Mean Square Error (MSE) between the reconstructed shapes and ground truth scans. The last row shows our result using the sample selected by the ranking network.
@@
Fig. 7 presents the qualitative results of 3D reconstructions from the generated shape coefficients by the network, alongside a comparison with the state-of-the-art neutral face reconstruction method MICA [55]. For quantitative evaluation, we utilize the NoW benchmark dataset [45], with the results presented in Tab. 1. The table demonstrates that our method achieves a lower average error across all reconstructed samples compared to other state-of-the-art occlusion-based reconstruction methods.
Figure 7: Neutral face reconstruction on occluded images. For the given occluded input image (a), (b) shows the reconstructed shape provided by MICA [55]; (c) is one of the generated samples from our method; (d) and (e) are the best and lowest-ranked samples, respectively, as selected by the ranking network.
5.3 Identity Ranking Network Evaluation
Experimental setup.
We utilized the same four datasets employed in IdGen. The network is trained for 120 epochs with a batch size of 32 on a Titan X GPU. The training process ran for 20 epochs per day, converging over the course of 6 days.
Baselines and Metrics.
We compare our approach with state-of-the-art occlusion-based reconstruction methods [14, 31]. We use the FLAME random sampling as a benchmark similar to IdGen. Our evaluation metrics include the mean, median, and standard deviation of the mean square error (MSE) as provided by the NoW benchmark [45]. A comprehensive comparison with other neutral face reconstruction methods [55, 52, 22] is also included in the supplementary material.
Results.
Fig. 7 illustrates the differences in reconstructed samples as identified by the ranking network. The variations in neutral face reconstructions are generally subtle, making it challenging to distinguish the best and worst samples qualitatively. To highlight the effectiveness of ranking in selecting high-quality samples, we conducted an ablation study, with results presented in Fig. 10. For quantitative evaluation, inspired by the approach in [31], we split our assessment into occluded versus non-occluded image reconstruction. The comparison results with other occlusion-based reconstruction methods on the NoW validation set, shown in Tab. 2, demonstrate that our model significantly outperforms the baselines.
Method Unoccluded subset Occluded subset
FLAME (Baseline) 1.79 2.24 1.89 1.80 2.26 1.91
Deep3D [13] 1.33 1.67 1.41 1.40 1.73 1.41
DECA [22] 1.18 1.47 1.24 1.29 1.56 1.29
MoFA [47] 1.35 1.69 1.42 1.36 1.69 1.41
FOCUS [31] 1.03 1.25 1.03 1.07 1.34 1.19
FOCUS (MP) [31] 1.02 1.24 1.02 1.08 1.34 1.20
OFER (selected by rank) 0.97 1.20 1.00 1.01 1.26 1.05
Table 2: Ranked sample evaluation in NOW validation benchmark. FLAME (baseline) is created by sampling random FLAME shape coefficients. We split the validation set into occluded and non-occluded subsets inspired by  [31] and evaluate against occlusion reconstruction methods using the median (Med), mean, and standard deviation (std) of the non-metrical Mean Square Error (MSE) between the reconstructed shapes and ground truth scans.
Figure 8: Sample ranking on NoW validation images. For image in column (a), we generate 100 shape coefficients with IdGen and rank them using IdRank. Column (b) shows the top-ranked sample, (c) a mid-ranked sample (50th), and (d) the lowest-ranked sample, based on median MSE against the ground truth scan. Columns (b)-(d) show viewpoints highlighting the sample errors, with median MSE shown beneath each sample.
5.4 Expression Generative Network Evaluation
Experimental setup.
We train the network using the FaMoS [5] dataset, specifically on a subset that captures diverse expressions across the entire dataset. The network was trained for 400 epochs with a batch size of 128 on an RTX 8000 GPU, with the training spanning one GPU day.
Figure 9: Comparison of expression reconstruction for in-the-wild occluded images. We compare against EMOCA [10] (showing front and side view, pink), two reconstructions from Diverse3D [14] (blue), and three samples (front and side view) from our method (green).
5.5 Final Reconstruction Evaluation
Our final reconstructed samples are obtained as explained in Sec. 4.4 and evaluated using the CO-545 validation dataset, which comprises of 545 frontal face images (no profile or rotated head views) synthetically occluded (hand, face, and random) across 11 subjects from COMA [41]. Furthermore, we assessed our reconstruction using the dataset provided by Dey et al. [14], which consists of 75 occluded images (mask, glasses, and random). Although this dataset does not differentiate between occluded and non-occluded regions, leading to less precise benchmarking, we included it for a comprehensive comparison.
Method $N_{C}^{unocc}$ $N_{C}^{occ}$ $E_{D}^{occ}$ $E_{C}^{unocc}$ $E_{C}^{occ}$
Diverse3D[14] 2.44 2.16 6.4 7.46 4.72
MICA[55] 1.75 1.83 NA NA NA
OFER (ours) 1.90 1.95 3.2 3.48 2.97
Table 3: Quantitative evaluation using RMS error metric for occluded image reconstruction. The letters represent: $N$ for Neutral, $E$ for expression, $Occ$ for occlusion, $unocc$ for no occlusion, $D$ for the dataset from dey et al. [14], and $C$ for the CO-545 dataset.
Baselines and Metrics.
We compare our method against the state-of-the-art diverse expression reconstruction approach by Dey et al. [14] and the single-hypothesis expression reconstruction method EMOCA [10]. It is important to note that Dey et al. [14] employs test-time optimization, whereas our approach does not. Additionally, their network is trained on the COMA dataset, which was used to create CO-545, making the evaluation set part of the training set. Our evaluation metric is the Root Mean Square Error (RMSE) between the unoccluded landmark vertices of the reconstructed samples and the ground truth.
Results.
The qualitative comparison of diverse expressive faces for occluded images is illustrated in Fig. 2 and Fig. 9. Our approach effectively generates plausible faces with varied expressions, even under challenging occlusions such as face masks. In contrast, EMOCA [10] tends to produce degenerate faces in similar hard occlusion scenarios, while Dey et al. [14] often results in unrealistic and extreme reconstructions. The quantitative evaluation shown in Tab. 3 supports this, particularly highlighting the substantial improvement over Dey et al. [14]. Our method demonstrates a significant advantage, reliably reconstructing visible regions and producing varied expressions in occluded areas.
6 Limitations and Future work
However, this introduces performance limitations sometimes due to the scarcity of such data. For effective model performance, it must learn from large amounts of high-quality data to aid in good generalization. Since we were able to obtain only a subset of available 3D supervision dataset, our method faces challenge in accurately reconstructing negative expressions. Therefore, a promising direction for future research is to integrate both 2D and 3D supervision to improve the performance of the method.
When we refer to performance, we specifically measure it in terms of how closely the reconstructed model approximates the ground truth scan. Although our ranking network selects a sample that minimizes error compared to a randomly chosen one from the generated set, it does not guarantee the selection of the absolute best match, i.e., the sample with the lowest error and closest approximation to the ground truth. From this observation we identify two potential research directions for constructing an ideal ranking network. The first is to refine the ranking architecture itself. A second promising approach involves integrating the ranking process directly within the diffusion process. Integrating ranking as a feedback mechanism within the diffusion process could improve the quality of the generated samples, resulting in improved overall performance.
7 Conclusion
In this paper, we present OFER, a conditional diffusion-based method for generating multiple hypotheses of expressive faces from a single-view, in-the-wild occluded image. We employ two diffusion models to generate FLAME 3DMM shape and expression coefficients. To ensure a consistent geometric face shape for varied expression of a single identity, we introduce a probabilistic ranking method to select an optimal sample from the generated shape coefficients. By combining the statistical learning strengths and generative capabilities of diffusion models, along with the smooth face reconstruction provided by a parametric model, our method produces plausible 3D faces that accurately reflect the input image. OFER achieves state-of-the-art results in diverse expression reconstruction, both qualitatively and quantitatively, significantly outperforming existing occlusion-based reconstruction methods.
Supplementary Material
8 Ablation Study
8.1 Identity Generative Network
The representation used for the conditioning image provided to the diffusion network plays an important role in the reconstruction task. In Tab. 4 we show ablation results for the embedding used in Identity Generative Network. The results indicate that using only the ArcFace [11] embedding improves the error metric of this network, compared to employing FaRL [53] or a combination of both as conditions. This may be because ArcFace [11] is used to distinguish identity-defining features, while FaRL [53] is designed for downstream facial analysis tasks that require capturing facial subtleties.
8.2 Identity Ranking Network
Benefits of combining ranking with reconstruction.
We conducted experiments to validate the importance of ranking as a selection process, presenting the results in Tab. 5. Our evaluation focuses on two approaches: optimization of the shape search space and selection of high-quality samples.
Ranking improves search space: In row (a), we show the results of generating 200 and 1000 random FLAME samples, while row (d) presents the results from OFER. The “ideal lowest error” column corresponds to an ideal ranking. Even with an increased number of random hypotheses (1000 samples), finding a good reconstruction in this space remains less likely, as indicated by the higher error in the ideal ranking scenario (0.81 with OFER vs 0.90 with FLAME).
@@ 8.2 Identity Ranking Network
Ranking selects low-error samples: In row (b), we show the results of ranking a set consisting of $50\%$ random FLAME samples and $50\%$ OFER reconstructions; row (c) reflects the same for an $80/20$ split. This mimics a scenario where the samples to rank include both correct solutions and high-error ones. Here, the ranked results (1st and 2nd columns) are consistently lower than the average error (3rd column). It indicates that low-error, high-quality samples are being selected. This is further supported by qualitative evidence in Fig. 10.
Embedding Occluded subset UnOccluded subset
ArcFace [11] 1.03 1.26 1.06 1.01 1.22 1.01
ArcFace+FaRL 1.05 1.30 1.09 1.03 1.28 1.06
FaRL [53] 1.15 1.40 1.12 1.07 1.30 1.06
Table 4: Ablation on image embeddings. MSE error on the NoW validation dataset, using ArcFace [11] embedding, FaRL [53] embedding, or a combination of both, for the Identity Generative Network.
Figure 10: Qualitative and quantitative results for best and least ranked samples for row (b) in Tab. 5 conditioned on NoW validation images. The median MSE error between the ground truth scan and the reconstructed 3D shape is displayed below each rendering. These results demonstrate that ranking as a selection method enhances the quality of sample selection.

### v2

@@ 3.3 Identity Ranking Network (IdRank)
Our goal is to generate a set of 3D faces with diverse expressions that align with the input occluded image. While expressions can vary, we would like here to preserve a unique and consistent identity across all samples. This approach is motivated by three main observations. First, we noticed that shape coefficients display much less variability than expression coefficients, as images typically contain more cues for the expression-independent facial features. Second, we noticed in our experiments that, when computing the reconstruction error for the N generated samples, the minimum value is on-par to state-of-the-art methods (see supplementary material). Since identity is typically more defined than expression, a method capable of selecting this optimal sample would be advantageous. Finally, even in cases of severe occlusions where identity features may be less discernible, having a method that effectively filters out poorly generated samples is important for practical applications.
@@
Towards this end, we introduce IdRank, a network specifically designed to rank the outputs of IdGen in order to select those that best preserve the facial structure of the input image.
Ranking framework.
The overall architecture of IdRank is shown in Fig. 5. Given a pool of N candidate shapes generated by IdGen, the goal is to select the one closest in distance to the ground-truth neutral shape (available during training). To achieve this, we train a ranking network $R_{\theta^{\prime}}$ that goes from a list of $v$ face vertices (derived from the shape coefficients) to a scalar $p$ representing the probability that the sample matches the ground truth. We then apply the network $R_{\theta^{\prime}}$ over the $N$ samples, ranking each based on their scores and selecting the highest-scored sample.
Before passing the mesh to the network, we remove the vertices corresponding to the back of the head since they are rarely visible in the image, and their inclusion in the error computation might lead to spurious results.
The number of candidates N is an important hyper-parameter, as it defines a trade-off between computational complexity and reconstruction error –a large value of $N$ means that the actual ground-truth shape has more chances of being generated, but incurs in a significantly higher computational cost. We empirically choose here $N=100$ (see supplementary material).
Generating ground-truth data.
During training, we generate the ground-truth data for ranking in an online fashion as follows. Given an image $I$ together with its ground-truth FLAME mesh $M_{GT}\in\mathbb{R}^{5023}$, we first sample $N$ shape parameters $\{\beta_{i}\}_{i=1}^{N}$ using IdGen (with gradient back-propagation disabled). We then reconstruct the neutral face vertices using Eq. 1, yielding a set of mesh candidates =${\cal M}=\{M_{i}=M(\beta,\mathbf{0},\mathbf{0})\in\mathbb{R}^{5023}\}_{i=1}^{N}$. From the set ${\cal M}$ as well as the ground-truth mesh $M_{GT}$ we retain only the frontal vertices using a pre-computed mask, resulting in ${\cal M}^{\text{frontal}}=\{\hat{M}_{i}\in\mathbb{R}^{n^{\prime}}\}_{i=1}^{N}$ and $M^{\text{frontal}}_{GT}\in\mathbb{R}^{n^{\prime}}$ ($n^{\prime}<5023$). Finally, we compute the error between each $\hat{M}_{i}$ and $M^{\text{frontal}}_{GT}$ to produce the following set:
${\cal D}_{GT}=\{|\hat{M}_{i}-M^{\text{frontal}}_{GT}|\}_{i=1}^{N}.$ (3)
with ||$\cdot$|| ⋅ | the L1 norm.
We empirically observed that removing the mean $\mu_{\mathcal{M}}=\frac{1}{N}\sum_{i=1}^{N}\hat{M}_{i}$ from the set ${\cal M}^{\text{frontal}}$ reduces redundancy and helps convergence of the network. Hence, we transform ${\cal M}_{\text{frontal}}$ into a zero-centered set of residual meshes ${\cal M}_{\text{centered}}$ via:
${\cal M}_{\text{center}}=\{M^{\prime}_{i}=(\hat{M}_{i}-\mu_{\mathcal{M}})\}_{i=1}^{N}.$ (4)
The final training set for image $I$ is defined as the list of pairs ${}\mathbf{X}=[(\mu_{\mathcal{M}},M^{\prime}_{i})]_{i=1}^{N}$, obtained by sorting the values in ${\cal D}_{GT}$ in ascending order.
Learning to rank facial shapes.
The ranking network $R_{\theta^{\prime}}$ is a small MLP that takes as input a single pair $(\mu_{\mathcal{M}},M_{i}^{\prime})$ and is trained to predict a sample score $d_{i}$. It is conditioned on a 1024-dimensional feature vector $c_{R}=(c_{a}\|c_{f})$, formed by concatenating features from the ArcFace ($c_{a}$[10] and FaRL ($c_{f}$) encoders [57], designed for facial analysis tasks. The network is run $N$ times producing ${\cal D}=\{d_{1},\dots,d_{N}\}$ scores, which are then passed through a softmax operator to transform them into probabilities, $P=\mathrm{softmax}({\cal D})=\{p_{i}=\frac{e^{d_{i}}}{\sum_{j=1}^{N}e^{d_{j}}}\}_{i=1}^{N}$. The ground-truth distance values ${\cal D}_{GT}$ are also passed through a softmax operator, ${\cal D}_{GT}^{\prime}=\mathrm{softmax}({\cal D}_{GT})$. Finally we compute the cross-entropy loss between the two distributions $P$ and ${\cal D}_{GT}^{\prime}$:
$L_{{\cal D}_{GT}^{\prime},P}(\theta^{\prime})=-\sum_{i=1}^{N}{{\cal D}_{GT}^{\prime}}(i)\log(p_{i}).$ (5)
@@
During inference, we select the sample with the highest probability score as the optimal sample, denoted as $S_{R}$. While multiple samples may have equal probability scores, their ordering in such cases would be rather random. In that case, we choose the first-ranked sample from this list of equally high probability samples.
Figure 5: Identity Ranking Network. Given the N shape coefficients from IdGen, we reconstruct the neutral meshes using FLAME. Each mesh is passed through a 5-layer MLP to compute a score, conditioned on the input image using the ArcFace [10] and FaRL [57] encoders. The N scores are then converted into probabilities using softmax. The ranking order of the sorted scores is compared against the ranking of the sorted reconstructed errors, and the network is trained to match them.
@@ 3.4 Expression Generative Network (ExpGen)
The expression network $\epsilon_{\theta^{\prime\prime}}^{E}$ architecture (Fig. 4, bottom) closely follows that of IdGen, employing the conditional DDPM framework to generate the first 50 components of the FLAME expression coefficients. It is conditioned on a 1024-dimensional feature vector, similar to IdRank.
During sampling, the network initializes from unit Gaussian noise $\varepsilon_{T}\in\mathbb{R}^{50}$ and iteratively denoises towards the FLAME expression parameters, $e_{0}\in\mathbb{R}^{50}$. The parameters $\theta^{\prime\prime}$ are optimized by minimizing the loss function Eq. 6
$L_{\varepsilon}(\theta^{\prime\prime})=\mathbb{E}_{t}\Bigl{[}\mathbb{E}_{e_{0},\varepsilon}[|\epsilon^{E}_{\theta}(e_{t})-\varepsilon_{t}|]\Bigr{]}.$ (6)
@@ 4 Experiments and Evaluations
In this section, we present qualitative and quantitative evaluation results for OFER. We begin by introducing the datasets used for training and testing, including the proposed CO-545 dataset, which we specifically designed to evaluate performance on occluded images (Sec. 4.1). Next, we provide evaluation metrics and experimental results on IdGen and IdRank (Sec. 4.2), followed by a similar analysis for the final reconstructions with ExpGen (Sec. 4.3).
@@ 4.1 Datasets
Training data.
We train IdGen with the Stirling [23], FaceWarehouse [7], LYHM [8], and Florence [3] datasets, which contain 2D-3D ground-truth pairs of images and corresponding registered FLAME shape coefficients. The same four datasets are used to train IdRank. For ExpGen we use the FaMoS (Facial Motion across Subjects) [5] dataset, which offers paired image and FLAME 3D meshes of faces in motion. The shape and expression coefficients were provided by the FaMos authors.
The CO-545 dataset.
To evaluate reconstruction of expressive faces under occlusions, we propose a new dataset built on top of the CoMA dataset [42]. The CO-545 validation dataset comprises 545 frontal face images (no profile or rotated head views) synthetically occluded (hand, face, and random) across 11 subjects taken from COMA [42]. For each image we identify the sets of occluded and unoccluded vertices, and evaluate them using separate metrics. More details can be found in the supplementary material.
Testing data.
We evaluate the final expressive reconstructions using the dataset provided by Dey et al. [13] (“Dey Dataset”), as well as the CO-545 dataset. Additionally, we test on the real-world hard occlusion images dataset provided by Erakiotan et al. [20], which does not have any 3D ground truth data. We additionally evaluate the quality of neutral shape reconstructions using the occluded and unoccluded subsets of the NoW benchmark [47]. Note that we do not have access to the occluded/unoccluded split for the testing set, and hence provide results mainly for the validation set. Additional comparisons on the test set are included in the supplementary material.
4.2 Evaluation on IdGen and IdRank
We first evaluate the accuracy of the neutral mesh reconstructions provided by the IdGen network, as well as the ranked samples according to IdRank. We trained IdGen for 400 epochs with a batch size of 128 on an RTX 8000 GPU. The entire training process spans one GPU day. IdRank was trained for 120 epochs with a batch size of 32 on a Titan X GPU. The training process ran for 20 epochs per day, converging over the course of 6 days.
@@ 4.2 Evaluation on IdGen and IdRank
Since OFER is designed to provide multiple solutions, we establish a baseline by randomly sampling FLAME shape coefficients from its parametric space. We evaluate our approach on the NoW validation benchmark, which allows separate measurement for the occluded and unoccluded subsets of the dataset. We compare our method to standard reconstruction techniques as well as state-of-the-art occlusion-based approaches, Diverse3D [13] and FOCUS [31]. Our evaluation metrics include the mean, median, and standard deviation of the mean square error (MSE) relative to the ground-truth mesh.
@@ Results.
Fig. 6 and Fig. 7 present qualitative results of 3D shape reconstruction, alongside comparisons with the state-of-the-art neutral face reconstruction method MICA [59] and the occluded face reconstruction method by Dey et al. [13], Diverse3D. In Fig. 6 we observe that, despite mild occlusions, OFER can recover accurate shapes with distinctive features (e.g., cheeks in the first row and the head length in the second row). We show an example of extreme occlusion in Fig. 7, where we reconstruct two images of the same person wearing a face mask, alongside a minimally occluded reference image. We observe that OFER, and in particular the top-ranked sample, produces plausible results for the two masked images.
@@
For quantitative evaluation, we use the NoW benchmark dataset [47], with the results presented in Tab. 2. Our method achieves a lower average error across all reconstructed samples compared to other state-of-the-art occlusion-based approaches. In addition, the shapes generated by IdGen can achieve competitive results with standard reconstruction methods, as shown by the minimum reconstruction error obtained over all samples (OFER-min).
Ablation.
@@
To highlight the effectiveness of ranking in improving the search space and in selecting high-quality samples, we conducted an ablation study, with results presented  Tab. 1. In row (a), we show the results of generating 200 and 1000 random FLAME samples, while row (d) presents the results from OFER. The “ideal lowest error” column corresponds to an ideal ranking (i.e. the minimum error within the samples). Even with an increased number of random hypotheses (1000 samples), finding a good reconstruction in this space remains less likely, as indicated by the higher error in the ideal ranking scenario (0.81 with OFER vs 0.90 with FLAME), showcasing the improved search space when ranking is applied to IdGen. In row (b) we show the results of ranking a set consisting of $50\%$ random FLAME samples and $50\%$ OFER reconstructions; row (c) reflects the same for an $80/20$ split. This mimics a scenario where the samples to rank include both correct solutions and high-error ones. Here, the ranked results (1st and 2nd columns) are consistently lower than the average error (3rd column), highlighting its effectiveness in selecting low-error samples. This is further supported by qualitative evidence in Fig. 8.
We further demonstrate the network’s ability to identify subtle differences to select top-ranked sample with lower error than the lower ranked samples, as shown in the color coded error map in Fig. 10 in the supplement.
Figure 6: Neutral face reconstruction on occluded images. For (a) the given occluded input image, (b) shows the reconstructed shape provided by MICA [59]; (c) is one of the generated samples from our method; (d) and (e) are the best and worst-ranked samples, respectively, as selected by the ranking network.
Figure 7: Neutral face reconstruction on a single identity. For (a) the given input image, (b) shows the reconstructed shape provided by MICA [59]; (c) shows the reconstructed shape provided by Diverse3D [13]; (d) and (e) are the worst and best-ranked samples as selected by the ranking method.